OpenStudy (anonymous):

what is the simpliest form of the quotient 4 sgrt 810 / 4 sgrt 2

OpenStudy (anonymous):

sqrt810=sqrt81*sqrt10=9sqrt10 so: 4sqrt810/4sqrt2 =4*9sqrt10/4sqrt2 =36sqrt10/4sqrt2 =9sqrt10/sqrt2 =9sqrt10*sqrt2/(sqrt2*sqrt2) =9sqrt20/2 =9sqrt4*sqrt5/2 =9*2sqrt5/2 =9sqrt5 There was probably a faster way to do that...

OpenStudy (anonymous):

that is what I got but ny answer choices are 4 sgrt 162 3 4sgrt5 5 4sgrt3 3 sgrt 5

OpenStudy (anonymous):

They're wrong :P Also it's "sqrt" not "sgrt" btw

OpenStudy (anonymous):

sorry wrong key

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

thanks